ues the 45-45-90 right triangle Conjucture to find the lenths of each triangle
a=5,b=__, c=__

Answer by checkley71(8403)

A^2+B^2=C^2
5^2+5^2=C^2
25+25=C^2
C^2=50
C=SQRT50
C=7.07 ANSWER.